A nonionizing solid dissolved in water change the freezing point to -2.79°C calculate its concentration.


1
Expert's answer
2021-05-09T00:00:44-0400

Freezing point = 0 ºC – ΔTf

"-2.79 = 0 - ΔT_f \\

ΔT_f = 2.79 \\

ΔT_f = K_f \times m \\

2.79 = 1.86 \times m \\

m = \frac{2.79}{1.86} = 1.5 \;M"

Answer: 1.5 M